OpenStudy (anonymous):

At the age of 30, to save for retirement, you decide to deposit $60 at the end of each month into an IRA that pays 3.5% compounded monthly. How much will you have in the IRA when you retire at 65?

OpenStudy (anonymous):

You have to use the compound interest formula: A=P(1 + r/n)^(nt) P is the initial amount you're investing. R is the percent that you're compounding at. T is time, and in the case, 65-30=35. N is going to be 12 because you are compounding something monthly and there are 12 months in a year. A=60(1 + 0.035/12)^(12*35)=$203.89